A brick is 50 cm long,40 cm wide and 10 cm thick.It weighs 24newton and is lying on the ground .find the pressure exerted

Grade:8

5 Answers

Arun
6 years ago

Hello!

By definition, a pressure is the force acting perpendicular to the surface divided by the area on which this force is distributed.

In our case, the force is the weight of a brick and it is given (24 N). The area on which the weight is distributed is equal to the length (50 cm) multiplied by the width (40 cm), so it is equal to 2000 cm2. So the pressure is

24/2000 = 0.012 (N/cm2)

Aashu Mittal
6 years ago
Pressure=Force(N)/Area(m^2)So, Area=length*breadth (in meter)Therefore, Area=1/2*2/5=1/5Now ,using above formula, Pressure=24*1/5=4.8 Pascal....Answer
Aashu Mittal
6 years ago
There was something wrong in last step, Actually this is correct method: Pressure=24/1/5=120 Pascal..Answer
Rushil
6 years ago
Here in questionP=F/AreaF=24NAREA =T.S.A OF CUBOID=2 (L*B +B*H+H*L)=2 (50 *40 +40*10+50*10)=6800cm ^2We know 1cm^2=0.0001m^2Thus,6800cm ^2 = 0.68m^2P=F/AP=24/0.68P=35.29 PASCAL
Nitesh
6 years ago
We know that pressure = force/area i.e perpendicular force/ area on which force acts.Since area is 50×40/10000 m2.And force given is 24N .So p=24×10000/2000=120 Pa